# Calibrated Value-Aware Model Learning with Stochastic Environment Models

### v1

@@ 9 Conclusions
Our experiments further show that the calibration of the $(m,b)$-VAML losses is important for obtaining strong learning with stochastic environment models. In addition, while previous papers showed that IterVAML losses can be unstable in practice (Lovatto et al., 2020; Voelcker et al., 2022), we find that this can be overcome by adopting the latent model architecture used in Schrittwieser et al. (2020) and the auxiliary losses established by Li et al. (2023); Hansen et al. (2022). When combined with a suitable value learning procedure, IterVAML performs on par with or better than MuZero in continuous control tasks. Finally, while $(m,b)$-VAML losses have mostly been used with deterministic environments, our work enables the community to use stochastic models with a calibrated loss and shows the potential merits of this approach in a number of environments.
@@ Acknowledgments
We would like to thank the members of the Adaptive Agents Lab and the Toronto Intelligent Systems Lab who provided feedback on the ideas, the draft of this paper, and provided additional resources for running all experiments. We would also like to thank the anonymous reviewers for providing valuable feedback to improve this work. AMF acknowledges the funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program (2021-03701). Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. We acknowledge Amin Rakhsha for helping with the formal statement of Lemma 4.
Societal Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Proofs and mathematical clarifications
$\mathbb{E}_{x\sim p,y\sim q(\cdot|x)}\left[\left|f(x)-y\right|^{2}\right]=\mathbb{E}_{x\sim p,y\sim q(\cdot|x)}\left[\left|f(x)-f^{*}(x)+f^{*}(x)y\right|^{2}\right],$
@@ A.1 Main propositions: Section 3
$[\mathcal{T}^{\mathcal{P}^{\pi}}V](x)=r(x)+\gamma\mathbb{E}_{x^{(1)}\sim\mathcal{P}^{\pi}(\cdot|x)}[V(x^{(1)})]=r(x)+\gamma\mathbb{E}_{x^{(1)}\sim\hat{p}(\cdot|x)}[V(x^{(1)})]=[\mathcal{T}^{\hat{p}}V](x)$ (11)
@@ A.2.1 Helper Lemmata
For the following results, we will use the following notation. Let $\mathcal{X}$ be a discrete sample space. Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a random variable. Let $\mathcal{P}^{*}$ be the set of distributions for which both $\mathcal{P}^{*}=\operatorname*{Arg\,min}_{p^{\prime}}\mathrm{Var}_{p^{\prime}}[f(x)]$, and for all $p^{*}\in\mathcal{P}^{*}$ $\mathbb{E}_{p^{*}}\left[f(x)\right]=\mathbb{E}_{p}\left[f(x)\right]$. Let $k$ be an integer. Finally, let $g(\xi)=\mathbb{E}_{x\sim p}\left[\left(\mathbb{E}_{y\sim\xi}\left[f(y)\right]-f(x)\right)^{2}\right]+\frac{1}{k}\mathrm{Var}_{\xi}[f(x)].$
@@
Let $p$ be a probability distribution over $\mathcal{X}$ with $\mathbb{E}_{p}\left[f(x)\right]$, for which no $x$ exists so that $f(x)=\mathbb{E}_{p}\left[f(x)\right]$. This is an assumption both on $f$ and $p$.
@@
$g(\xi^{*})$ $=\mathbb{E}_{x\sim p}\left[\left(\mathbb{E}_{y\sim\xi^{*}}\left[f(y)\right]-f(x)\right)^{2}\right]+\frac{1}{k}\mathrm{Var}_{\xi}[f(x)]$ (12)
@@
$=\mathrm{Var}_{p}(f(x))+\frac{1}{k}\mathrm{Var}_{\xi^{*}}[f(x)]$ (14)
@@
$<\mathrm{Var}_{p}[f(x)]+\frac{1}{k}\mathrm{Var}_{p^{*}}[f(x)]=g(p^{*})$ (24)
@@ A.2.3 Main results: MuZero
By the Cauchy-Schwartz inequality, the variance is only 00 for a $g(x)$ constant almost everywhere. However, this violates the assumption. Therefore there exists an $\varepsilon>0$ so that $\mathcal{L}\left(\hat{g}\right)\leq\mathcal{L}\left(g\right)$. We now can construct the function space $\mathcal{F}$ so that it includes at least $g$ and $g+\varepsilon g.$
@@ A.3.1 Propositions from Bertsekas & Shreve (1978)
For convenience, we quote some results from Bertsekas & Shreve (1978). These are used in the proof of Lemma 5. While some of the definitions are rather technical, it is mostly sufficient to see a stochastic kernel as the continuous generalization of the transition matrix in finite MDPs. The projection used in 2 is simply a restriction of a set of tuples $(x,y)$ on the $x$ values. Other topological statements are standard and can be found in textbooks such as Munkres (2018).
@@ Appendix B Implementation Details – Section 5: Garnet
Each Garnet is defined by a randomly generated transition matrix and a random reward function. Across all experiments we use 50 dimensional state spaces. For each state $x_{i}$, we sample $k=10$ successor states at random without replacement. Then we assign each successor state $x_{j}$ a weight $\omega_{ij}\sim\mathcal{N}(0,1)$ and set the weight of all non-successor states to the floating point minimum. The transition matrix is computed via the softmax function. The reward is sampled as $r(x_{i})\sim\mathcal{N}(0,1).$
@@ C.1 Architecture
The original MuZero paper (Schrittwieser et al., 2020) provides large-scale experiments on the ALE benchmark (Bellemare et al., 2013). However, the original algorithm and the ALE suite are computationally extremely expensive to run. In addition, to the best of our knowledge, no reliable open-source replication of the closed-source MuZero and Stochastic MuZero models and algorithms exists. Instead, we use the TD-MPC architecture and algorithm (Hansen et al., 2022) as our main reference. This implementation uses the MuZero loss but is significantly cheaper and faster to run in a small-scale experimental setting, while still solving non-trivial control tasks.
@@
Code is provided in the supplementary files and open-sourced after the review process.
@@ C.3 Additional results
In addition to the raw performance graphs in the main paper, we present the average model entropy and VAML loss curves in Figure 5. As expected, we find that the CVAMl loss leads to a higher entropy model compared to the VAML loss. This however does not lead to a noticeable difference in the VAML error itself. We also present the individual reward curves for each task in both domains in Figure 6.
Figure 5: Model entropy (left) and loss value for the VAML-losses (right) aggregated across all environments. The model entropy of the corrected $(1,0)$-VAML differs significantly. However, this does not translate to a pronounced difference in the VAML error itself. Therefore, while the calibration term does lead to a higher variance, this does not necessarily translate into an improvement in performance. In $(1,1)$-VAML compared with the calibrated version, the model entropy does not differ significantly. However, without calibration, the VAML error seems to be growing slightly stronger. Given the large deviation in performance we observe in the main results for $(1,1)$-VAML, we conjecture that correcting the value function learning term is significantly more important for performance.
Figure 6: Episode rewards normalized by maximum reward per environment averaged over 30 seeds.
@@ C.5 Garnet Policy Iteration
Figure 9: Return curves for cliffwalk policy iteration. In each iteration, the policy’s value function is estimated with the model-based losses, following the garnet experiments in the main paper. The policy is improved in a greedy fashion. The problem is a modified 5x5 cliffwalk environment https://gymnasium.farama.org/environments/toy_text/cliff_walking/ where the likelihood of moving in the intended direction is given by temp. Each row is a different model rank (see description in the main paper), each column a different temperature. Results are consistent with the garnet experiments in the main paper, except for the surprising outlier of the uncorrected $(1,1)$ loss in the almost-deterministic case. We found that the dependency on the real reward can stabilize the performance, and the low model rank prevents the pathologies discussed in the paper. However, this only holds for low rank deterministic case. Each curve shows confidence intervals over 1000 seeds.

### v2

@@ 9 Conclusions
Our experiments further show that the calibration of the $(m,b)$-VAML losses is important for obtaining strong learning with stochastic environment models. In addition, while previous papers showed that IterVAML losses can be unstable in practice (Lovatto et al., 2020; Voelcker et al., 2022), we find that this can be overcome by adopting the latent model architecture used by Schrittwieser et al. (2020) and the auxiliary losses established by Li et al. (2023); Hansen et al. (2022). When combined with a suitable value learning procedure, IterVAML performs on par with or better than MuZero in continuous control tasks. Finally, while $(m,b)$-VAML losses have mostly been used with deterministic environments, our work enables the community to use stochastic models with a calibrated loss and shows the potential merits of this approach in a number of environments.
@@ Acknowledgments
We would like to thank the members of the Adaptive Agents Lab and the Toronto Intelligent Systems Lab who provided feedback on the ideas, the draft of this paper, and provided additional resources for running all experiments. We acknowledge Amin Rakhsha for helping with the formal statement of Lemma 4. We would also like to thank the anonymous reviewers for providing valuable feedback to improve this work. AMF acknowledges the funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program (2021-03701). Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Impact Statement
Our paper is first and foremost a theoretical analysis of loss function used in reinforcement learning. As such, it presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Proofs and mathematical clarifications
$\mathbb{E}_{x\sim p,y\sim q(\cdot|x)}\left[\left|f(x)-y\right|^{2}\right]=\mathbb{E}_{x\sim p,y\sim q(\cdot|x)}\left[\left|f(x)-f^{*}(x)+f^{*}(x)-y\right|^{2}\right],$
@@ A.1 Main propositions: Section 3
$[\mathcal{T}^{\mathcal{P}^{\pi}}V](x)=r(x)+\gamma\mathbb{E}_{x^{(1)}\sim\mathcal{P}^{\pi}(\cdot|x)}[V(x^{(1)})]=r(x)+\gamma\mathbb{E}_{x^{(1)}\sim\hat{p}(\cdot|x)}[V(x^{(1)})]=[\mathcal{T}^{\hat{p}}V](x).$ (11)
@@ A.2.1 Helper Lemmata
For the following results, we will use the following notation. Let $\mathcal{X}$ be a discrete sample space and $p$ a distribution over it. Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a random variable. Let $\mathcal{P}^{*}$ be the set of distributions for which both $\mathcal{P}^{*}=\operatorname*{Arg\,min}_{p^{\prime}}\mathrm{Var}_{p^{\prime}}[f(x)]$, and for all $p^{*}\in\mathcal{P}^{*}$ $\mathbb{E}_{p^{*}}\left[f(x)\right]=\mathbb{E}_{p}\left[f(x)\right]$. Let $k$ be an integer. Finally, let $g(\xi)=\mathbb{E}_{x\sim p}\left[\left(\mathbb{E}_{y\sim\xi}\left[f(y)\right]-f(x)\right)^{2}\right]+\frac{1}{k}\mathrm{Var}_{\xi}[f(x)].$
@@
Let $p$ be a probability distribution over $\mathcal{X}$ with $\mathbb{E}_{p}\left[f(x)\right]$, for which no $x$ exists so that $f(x)=\mathbb{E}_{p}\left[f(x)\right]$. This is an assumption on both $f$ and $p$.
@@
$g(\xi^{*})$ $=\mathbb{E}_{x\sim p}\left[\left(\mathbb{E}_{y\sim\xi^{*}}\left[f(y)\right]-f(x)\right)^{2}\right]+\frac{1}{k}\mathrm{Var}_{\xi^{*}}[f(x)]$ (12)
@@
$=\mathrm{Var}_{p}(f(x))+\frac{1}{k}\mathrm{Var}_{\xi^{*}}[f(x)].$ (14)
@@
$<\mathrm{Var}_{p}[f(x)]+\frac{1}{k}\mathrm{Var}_{p^{*}}[f(x)]=g(p^{*}).$ (24)
@@ A.2.3 Main results: MuZero
By the Cauchy-Schwarz inequality, the variance is only 00 for a $g(x)$ constant almost everywhere. However, this violates the assumption. Therefore there exists an $\varepsilon>0$ so that $\mathcal{L}\left(\hat{g}\right)\leq\mathcal{L}\left(g\right)$. We now can construct the function space $\mathcal{F}$ so that it includes at least $g$ and $g+\varepsilon g.$
@@ A.3.1 Propositions from Bertsekas & Shreve (1978)
For convenience, we quote some results from Bertsekas & Shreve (1978). These are used in the proof of Lemma 5. While some of the definitions are rather technical, it is mostly sufficient to see a stochastic kernel as the continuous generalization of the transition matrix in finite MDPs. The projection used in 2 is simply a restriction of a set of tuples $(x,y)$ on the $x$ values. Other topological statements are standard and can be found in textbooks such as Munkres (2018). In the following $\mathcal{C}$ refers to sets of continuous functions.
@@ Appendix B Implementation Details – Section 5: Garnet
Each Garnet is defined by a randomly generated transition matrix and a random reward function. Across all experiments we use state spaces of size $|\mathcal{X}|=50$. For each state $x_{i}$, we sample $k=10$ successor states at random without replacement. Then we assign each successor state $x_{j}$ a weight $\omega_{ij}\sim\mathcal{N}(0,1)$ and set the weight of all non-successor states to the floating point minimum. The transition matrix is computed via the softmax function. The reward is sampled as $r(x_{i})\sim\mathcal{N}(0,1).$
@@ C.1 Architecture
The original MuZero paper (Schrittwieser et al., 2020) provides large-scale experiments on the ALE benchmark (Bellemare et al., 2013). However, the original algorithm and the ALE suite are computationally extremely expensive to run. In addition, to the best of our knowledge, no reliable open-source replication of the closed-source MuZero and Probabilistic MuZero models and algorithms exists. Instead, we use the TD-MPC architecture and algorithm (Hansen et al., 2022) as our main reference. This implementation uses the MuZero loss but is significantly cheaper and faster to run in a small-scale experimental setting, while still solving non-trivial control tasks.
@@ C.3 Additional results
In addition to the raw performance graphs in the main paper, we present the average model entropy and VAML loss curves in Figure 5. As expected, we find that the CVAML loss leads to a higher entropy model compared to the VAML loss. We compute the empirical differential entropy of the model prediction over the replay buffer as
$-\sum_{x,a\in\mathcal{D}}p(z^{\prime}|\varphi(x),a)\log p(z^{\prime}|\varphi(x),a),$
with $z^{\prime}$ sampled from the latent model. This however does not lead to a noticeable difference in the VAML error itself. We also present the individual reward curves for each task in both domains in Figure 6.
Figure 5: Model entropy (left) and loss value for the VAML-losses (right) aggregated across all environments. The model entropy of $(1,0)$-CVAML differs significantly. However, this does not translate to a pronounced difference in the VAML error itself. Therefore, while the calibration term does lead to a higher variance, this does not necessarily translate into an improvement in performance. In $(1,1)$-VAML compared with the calibrated version, the model entropy does not differ significantly. However, without calibration, the VAML error seems to be growing slightly stronger. Given the large deviation in performance we observe in the main results for $(1,1)$-VAML, we conjecture that correcting the value function learning term is significantly more important for performance.
Figure 6: Episode returns per environment averaged over 30 seeds.
@@ C.4 Ablations
We provide additional ablations on the model loss here.
In the first set of experiments (Figure 7), we remove the CVAML loss from the model. We observe deteriorating agent performance, especially with the stochastic models, and in the humanoid tasks. In addition, we compare against a model-free TD3 baseline that uses the same architecture as our model based versions.
In the second set (Figure 8), we ablate the auxiliary loss. Again, we find that the auxiliary loss helps especially in the humanoid experiments. In addition, the $(1,1)$-CVAML baseline suffers significantly more from the removal of the auxiliary loss than the $(1,0)$-CVAML loss. Upon further investigation we observe that the bootstrapped TD target of the $(1,1)$-CVAML loss leads to less stable learning compared to the value function target.
@@ C.5 Garnet Policy Iteration
Here we extend the Garnet experiments in Section 5 to the control case. In each iteration, the policy’s value function is estimated with the model-based losses, following the Garnet experiments in the main paper. The policy is improved in a greedy fashion. The problem is a modified 5x5 cliffwalk environment https://gymnasium.farama.org/environments/toy_text/cliff_walking/ where the likelihood of moving in the intended direction is given by temp. Results are presented in Figure 9
Results are consistent with the garnet experiments in the main paper, except for the surprising outlier of the uncalibrated $(1,1)$-VAML loss in the almost-deterministic case. This might be due to the different structures of the cliffwalk environment compared to the more diverse Garnet environments. However, this only holds for low rank deterministic case (top right corner).
Figure 9: Return curves for cliffwalk policy iteration. Each curve shows confidence intervals over 1000 seeds. Each row is a different model rank (see description in the main paper), each column a different temperature.